The Color Box uses a combination of common stock, preferred stock, and debt financing. The company wants preferred stock to represent 7 percent of the total financing. It also wants to structure the firm in a manner that will produce a weighted average cost of capital of 9.5 percent. The aftertax cost of debt is 4.8 percent, the cost of preferred is 8.9 percent, and the cost of common stock is 14.7 percent. What percentage of the firm's capital funding should be debt financing?

  1. 48.42 percent
  2. 52.03 percent
  3. 54.15 percent
  4. 44.78 percent
  5. 39.21 percent
Answer #1

WACC = WEIGHTED AVERAGE COST OF CAPITAL.  

It's the overall cost of raising funds for the firm.

Weighted average cost of capital= (weight of equity x cost of equity)+(weight of debt x cost of debt)+(weight of preference x cost of preference)

Weight of equity (we) = 1-weight of preference-weight of debt
we = 1-0.07-wd
we= 0.93-wd

Answer: wd = 48.42%

wack = weke & wd kd & wpkp walc = (0-93 - wd) Ice + wdled two plep 9-5% - (0.93-wd 1407 + ( wdx408)+ (0.07 x 8.9%) 9.5 = 13.6